# Some Primal-Dual Theory for Subgradient Methods for Strongly Convex Optimization

### v3

@@ 3.1 Statement of Primal-Dual Convergence Guarantees
Theorem 3.2 further recovers and extends the classic linear convergence of proximal gradient descent for smooth, strongly convex optimization. Assume $m=0$ and $f_{0}$ is $\beta$-smooth with $g_{0}(x;\xi)=\nabla f_{0}(x)$. Then (2.8) holds with $L_{0}=0$ and $L_{1}=2\beta$ [^4] This can be verified by noting $h_{x_{\mathtt{OPT}}}$ is $\beta$-smooth. Then the standard descent lemma ensures $h_{x_{\mathtt{OPT}}}(x_{\mathtt{OPT}})\leq h_{x_{\mathtt{OPT}}}(x_{k}-(\nabla f_{0}(x_{k})+n_{x_{\mathtt{OPT}}})/\beta)\leq h_{x_{\mathtt{OPT}}}(x_{k})-\frac{1}{2\beta}\|\nabla f_{0}(x_{k})+n_{x_{\mathtt{OPT}}}\|^{2}_{2}.$ Consider the stepsize selection with $\alpha_{0}=1/\mu$ and $\alpha_{k}=1/L_{1}$ constant thereafter, which corresponds to dual weights $\lambda_{0}=1$ and $\lambda_{k}=(1-\mu/L_{1})^{-(k-1)}/(L_{1}/\mu-1)$. This choice has $T_{0}=0$ and $C_{0}=(\frac{L_{1}}{\mu}-1)\delta_{0}(x_{\mathtt{OPT}})$, giving the following linear convergence
$\mathtt{primal\mbox{-}gap}_{T}+\mathtt{dual\mbox{-}gap}_{T}+\frac{\mu}{2}\mathbb{E}_{\xi}\|x_{T}-x_{\mathtt{OPT}}\|_{2}^{2}\leq\frac{L_{1}}{\mu}\delta_{0}(x_{\mathtt{OPT}})\left(1-\frac{\mu}{L_{1}}\right)^{T}$
@@
if $C_{0}$ is finite and $\sum_{k=0}^{T}\lambda_{k}\alpha_{k}/\sum_{k=0}^{T}\lambda_{k}\rightarrow 0$. The classic conditions needed for limiting primal convergence under Lipschitz continuity are $\alpha_{k}\rightarrow 0$, which implies $C_{0}$ is finite, and $\sum^{T}_{k=0}\alpha_{k}^{2}/\sum^{T}_{k=0}\alpha_{k}\rightarrow 0$, which differs slightly from our theory when $\alpha_{k}=\lambda_{k}/\mu\sum_{i=0}^{k}\lambda_{i}$ as
@@
One can select stepsizes to minimize our rate. Given $C_{0}=0$ and the first $T$ stepsizes $\alpha_{0},\dots\alpha_{T-1}$ and corresponding weights $\lambda_{0},\dots\lambda_{T-1}$, one can select $\alpha_{T}$ and $\lambda_{T}$ to minimize our convergence upper bound (1.5) after one more step by setting
@@
Without strong convexity, one cannot guarantee convergence of a duality gap since a linear $M^{(k)}$ leads the duality gap to always be 00 or $\infty$. Our theory can still be applied by a standard trick: Instead of unconstrained minimization of a convex function $f_{0}$, one could minimize the closely related strongly convex function
@@
This perturbed problem has minimum value at most $p_{\star}+\epsilon\frac{\|x^{*}-x_{0}\|^{2}_{2}}{2D^{2}}$, and so any $\epsilon$-minimizer of $\tilde{f}_{0}$ is an $(1+\frac{\|x^{*}-x_{0}\|^{2}_{2}}{2D^{2}})\epsilon$-minimizer for the original problem.
Note $\tilde{f}_{0}$ is $\epsilon/D^{2}$-strongly convex and if $f_{0}$ was $M$-Lipschitz continuous, then the perturbed objective $\tilde{f}_{0}$ satisfies (2.8) with $L^{2}_{0}=8M^{2}$ and $L_{1}=4\epsilon/D^{2}$. As a result, Theorem 3.2 ensures applying the subgradient method to $\tilde{f}_{0}$ with stepsize $\alpha_{k}=2D^{2}/\epsilon(k+2)$ has perturbed primal-dual gap converge at a rate $\frac{32M^{2}D^{2}}{\epsilon(T+1)}+O(1/T^{2})$. Similar perturbed primal-dual guarantees using a novel proof method were given by [27].
@@ 3.2 Proof of Primal-Dual Convergence Guarantees
This proof follows a standard analysis technique, directly expanding the definition of $R_{k+1}^{2}(y)$. First, suppose $x_{k}$ is feasible. Then
@@
where the first line uses that $\mathrm{prox}_{\alpha_{k},r}(y+\alpha_{k}n_{y})=y$, the second uses the nonexpansiveness of the proximal operator [34, Proposition 12.19], the third factors the norm squared and uses $\alpha_{k}=\lambda_{k}/w_{k}$, and the fourth uses the strong convexity of $h_{y}$. Then, applying the bound (2.8) gives the claim. Similarly, supposing $x_{k}$ is infeasible gives
@@
$D_{k}:=\left(\sum_{i<k:s(x_{i})=0}\lambda_{i}\right)p_{\star}-\mathbb{E}_{\xi}\inf M^{(k-1)}\ .$ (3.4)
Dividing through by $\sum_{i<k:s(x_{i})=0}\lambda_{i}$, once nonzero, gives the dual gap.
@@
$D_{k+1}\leq D_{k}-\frac{\lambda_{k}}{2}\left(\left(2-L_{1}\alpha_{k}\right)\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}})-L_{0}^{2}\alpha_{k}\right)\ .$
@@
Observe that one can rewrite $M^{(k)}(x)=Q_{1}(x)+Q_{2}(x)$ as the sum of two quadratics where $Q_{1}(x)=M^{(k-1)}(x)$ and $Q_{2}$ depends on whether $x_{k}$ is feasible. In the notation of Lemma 2.2, Lemma 2.1 ensures $Q_{1}$ has $a_{1}=\inf M^{(k-1)}$, $b_{1}=\mu\sum_{i=0}^{k-1}\lambda_{i}$, and $z_{1}=x_{k}$. To determine $Q_{2}$, first suppose $x_{k}$ is feasible. Then we have
$Q_{2}(x)=f_{0}(x_{k})+\langle g_{k},x-x_{k}\rangle+\frac{\mu}{2}\|x-x_{k}\|_{2}^{2}+r(x_{k+1})+\langle n_{k+1},x-x_{k+1}\rangle\ .$
@@
$a_{2}$ $=f_{0}(x_{k})+r(x_{k+1})+\alpha_{k}\langle n_{k+1},g_{k}+n_{k+1}\rangle-\frac{1}{2\mu}\|g_{k}+n_{k+1}\|_{2}^{2}\ ,$
@@
Lower bounding $r(x_{k+1})$ by $r(x_{\mathtt{OPT}})+\langle n_{x_{\mathtt{OPT}}},x_{k+1}-x_{k}+x_{k}-x_{\mathtt{OPT}}\rangle$, this simplifies to
$\mathbb{E}_{\xi}\inf M^{(k)}\geq\mathbb{E}_{\xi}\inf M^{(k-1)}+\lambda_{k}\left(\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}})+p_{\star}-\frac{\alpha_{k}}{2}\mathbb{E}_{\xi}\|g_{k}+n_{x_{\mathtt{OPT}}}\|_{2}^{2}\right)\ .$
In terms of $D_{k}=\sum_{i<k:s(x_{i})=0}\lambda_{i}p_{*}-\mathbb{E}_{\xi}\inf M^{(k-1)}$, this gives the following recurrence
@@
Applying (2.8) gives the claim in this case. Now suppose $x_{k}$ is infeasible. Then, noting $Q_{2}$ minimizes at $x_{k}-\frac{g_{k}}{\mu}$, one has
@@
$\mathbb{E}_{\xi}\inf M^{(k-1)}+\lambda_{k}\mathbb{E}_{\xi}\left(f_{s(x_{k})}(x_{k})-\frac{1}{2\mu}\|g_{k}\|_{2}^{2}\right)+\frac{\lambda_{k}\sum_{i=0}^{k-1}\lambda_{i}}{\mu\sum_{i=0}^{k}\lambda_{i}}\mathbb{E}_{\xi}\left\|g_{k}\right\|_{2}^{2}$
@@
In terms of $D_{k}=\sum_{i<k:s(x_{i})=0}\lambda_{i}p_{*}-\mathbb{E}_{\xi}\inf M^{(k-1)}$, this gives the recurrence
$D_{k+1}\leq D_{k}-\lambda_{k}\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}})+\frac{\lambda_{k}\alpha_{k}}{2}\mathbb{E}_{\xi}\|g_{k}\|_{2}^{2}\ .$
Bounding $\mathbb{E}_{\xi}\|g_{k}\|_{2}^{2}$ by (2.8) gives the claim in this last case. ∎
@@
$\mathbb{E}_{\xi}\|x_{T}-y\|^{2}_{2}$ $\leq\left(1+\frac{\max\{2,L_{1}/\mu-2\}L_{1}}{\mu}\right)^{T}\left(\|x_{0}-y\|^{2}_{2}+\frac{L_{0}^{2}}{L_{1}}+\frac{L_{0}^{2}}{\mu\max\{2,L_{1}/\mu-2\}}\right)\ .$
@@
The first inequality lower bounding $\delta_{k}(y)$ is immediate from (2.8). For the second inequality, note that if $x_{k}$ is feasible, convexity of $h_{y}$ ensures that $h_{y}(y)\geq h_{y}(x_{k})+\langle\mathbb{E}_{\xi_{k}}(g_{s(x_{k})}(x_{k};\xi_{k})+n_{y}),y-x_{k}\rangle$. If $x_{k}$ is infeasible, $f_{s(x_{k})}(y)\geq f_{s(x_{k})}(x_{k})+\langle\mathbb{E}_{\xi_{k}}g_{s(x_{k})}(x_{k};\xi_{k}),y-x_{k}\rangle$. In either case, Cauchy-Schwarz and Assumption C give the second inequality. This second bound implies $\delta_{k}(y)\leq\frac{L_{0}^{2}}{L_{1}}+L_{1}\|x_{k}-y\|_{2}^{2}$, completing the proposition’s first claim. To prove the proposition’s second claim, note that
@@
$\leq\left(1+\frac{\max\{2,L_{1}/\mu-2\}L_{1}}{\mu}\right)\mathbb{E}_{\xi}\|x_{k}-y\|^{2}_{2}+\frac{\max\{2,L_{1}/\mu-2\}}{\mu}\frac{L_{0}^{2}}{L_{1}}+\frac{L_{0}^{2}}{\mu^{2}}$
where the first inequality uses Lemma 3.2 divided by $(\frac{\mu}{2}\sum_{i=0}^{k}\lambda_{i})$, the second inequality applies simple upper bounds, the third inequality uses our bound on $|\delta_{k}(y)|$, and the fourth uses that $\alpha_{k}\leq 1/\mu$. From this, the proposition’s second claim follows as such recurrences of the form $a_{k+1}\leq b\cdot a_{k}+c$ satisfy $a_{T}\leq b^{T}a_{0}+c\frac{b^{T}-1}{b-1}\leq b^{T}(a_{0}+\frac{c}{b-1})$. ∎

### v5

@@ 3.1 Statement of Primal-Dual Convergence Guarantees
Theorem 3.2 further recovers and extends the classic linear convergence of proximal gradient descent for smooth, strongly convex optimization. Assume $m=0$ and $f_{0}$ is $\beta$-smooth with $g_{0}(x;\xi)=\nabla f_{0}(x)$. Then (2.8) holds with $L_{0}=0$ and $L_{1}=2\beta>\mu$ [^3] This can be verified by noting $h_{x_{\mathtt{OPT}}}$ is $\beta$-smooth. Then the standard descent lemma ensures $h_{x_{\mathtt{OPT}}}(x_{\mathtt{OPT}})\leq h_{x_{\mathtt{OPT}}}(x_{k}-(\nabla f_{0}(x_{k})+n_{x_{\mathtt{OPT}}})/\beta)\leq h_{x_{\mathtt{OPT}}}(x_{k})-\frac{1}{2\beta}\|\nabla f_{0}(x_{k})+n_{x_{\mathtt{OPT}}}\|^{2}_{2}.$ Consider the stepsize selection with $\alpha_{0}=1/\mu$ and $\alpha_{k}=1/L_{1}$ constant thereafter, which corresponds to dual weights $\lambda_{0}=1$ and $\lambda_{k}=\frac{\mu}{L_{1}}(1-\mu/L_{1})^{-k}$ , for $k>0$ This choice has $T_{0}=0$ and $C_{0}=(\frac{L_{1}}{\mu}-1)\delta_{0}(x_{\mathtt{OPT}})$, giving the following linear convergence
$\mathtt{primal\mbox{-}gap}_{T}+\mathtt{dual\mbox{-}gap}_{T}+\frac{\mu}{2}\|x_{T}-x_{\mathtt{OPT}}\|_{2}^{2}\leq\frac{L_{1}}{\mu}\delta_{0}(x_{\mathtt{OPT}})\left(1-\frac{\mu}{L_{1}}\right)^{T},$
@@
if $C_{0}$ is finite and $\sum_{k=0}^{T}\lambda_{k}\alpha_{k}/\sum_{k=0}^{T}\lambda_{k}\rightarrow 0$. Note this implies $\sum_{k=0}^{T}\lambda_{k}\rightarrow\infty$ since $\alpha_{k},\lambda_{k}>0$. The classic conditions needed for limiting primal convergence under Lipschitz continuity are $\alpha_{k}\rightarrow 0$, which implies $C_{0}$ is finite, and $\sum^{T}_{k=0}\alpha_{k}^{2}/\sum^{T}_{k=0}\alpha_{k}\rightarrow 0$, which differs slightly from our theory when $\alpha_{k}=\lambda_{k}/\mu\sum_{i=0}^{k}\lambda_{i}$ as
@@
One can select stepsizes to minimize our rate. Given $C_{0}=0$ and the first $T$ stepsizes $\alpha_{0},\dots\alpha_{T-1}$ and corresponding weights $\lambda_{0},\dots\lambda_{T-1}$, one can select $\alpha_{T}$ and $\lambda_{T}$ to minimize our convergence upper bound (1.5) after one more step by setting [^4] The formula for the optimal $\lambda_{T}$ in (3.5) can be verified as the unique solution to $\frac{d}{d\lambda_{T}}\left(\frac{\sum_{k=0}^{T}\lambda_{k}\alpha_{k}}{\sum_{k=0}^{T}\lambda_{k}}\right)=0$.
@@
For deterministic settings without regularization (i.e., supposing $g_{s}(x;\xi)$ is independent of $\xi$ and $r=0$ which allow expectations and $n_{y}$ to be omitted from Theorem 3.2 respectively) where $\mu$ and an upper bound $G^{2}\geq L_{0}^{2}$ are known, one can also utilize our theory to adapt stepsizes to avoid any early exponential divergences. Recall such divergences are quantified by $C_{0}$ as discussed at the end of Section 2. If one selects decreasing stepsizes with $\alpha_{0}=1/\mu$ and $\alpha_{k}\leq 1/L_{1}$ thereafter and $\lambda_{0}=1$, $T_{0}$ defined in (2.9) is at most zero, and hence no bad divergence can occur as
$\left(\frac{\sum_{k<T:s(x_{k})=0}\lambda_{k}}{\sum_{k=0}^{T-1}\lambda_{k}}\right)\left(\mathtt{primal\mbox{-}gap}_{T}+\mathtt{dual\mbox{-}gap}_{T}\right)$
$\qquad\qquad\leq\frac{G^{2}\sum_{k=0}^{T-1}\lambda_{k}\alpha_{k}+(\frac{1}{\alpha_{1}\mu}-1)\frac{\|g_{s(x_{0})}(x_{0};\xi_{0})\|^{2}}{2\mu}}{\sum_{k=0}^{T-1}\lambda_{k}}$ (3.6)
where we bounded $L_{0}^{2}\leq G^{2}$ and $C_{0}\leq(\frac{L_{1}}{\mu}-1)\delta_{0}(y)\leq(\frac{1}{\alpha_{1}\mu}-1)\frac{\|g_{s(x_{0})}(x_{0};\xi_{0})\|^{2}}{2\mu}$555 The inequality $\delta_{0}(y)\leq\|g_{s(x_{0})}(x_{0};\xi_{0})\|^{2}/2\mu$ follows from the strong convexity of $f_{s(x_{0})}$ and $n_{y}=0$ as $\delta_{0}(y)$ $=f_{s(x_{0})}(x_{0})-f_{s(x_{0})}(y)$ $\leq f_{s(x_{0})}(x_{0})-\inf_{z}f_{s(x_{0})}(z)$ $\leq f_{s(x_{0})}(x_{0})-\inf_{z}\left(f_{s(x_{0})}(x_{0})+\langle g_{s(x_{0})}(x_{0};\xi_{0}),z-x_{0}\rangle+\frac{\mu}{2}\|z-x_{0}\|^{2}_{2}\right)$ $=\|g_{s(x_{0})}(x_{0};\xi_{0})\|^{2}/2\mu.$ for either $y\in\{x_{\mathtt{OPT}},x_{\mathtt{SL}}\}$. Notice every quantity in (3.6) is computable! Hence, if one selected generic decreasing stepsizes $\alpha_{k}$, without knowing $L_{1}$ to ensure $\alpha_{k}\leq 1/L_{1}$ for $k\geq 1$, one can still check if convergence is occurring at the above rate. If (3.6) fails at some iteration, one can conclude $\alpha_{1}>1/L_{1}$. In this case, one could reasonably restart the method with reduced stepsizes, via an exponential backtracking.
Remark 8.
Without strong convexity, one cannot guarantee convergence of a duality gap since a linear $M^{(k)}$ leads the duality gap to always be 00 or $\infty$. Our theory can still be applied by a standard trick: Instead of unconstrained minimization ($m=0$) of a convex function $f_{0}$, one could minimize the closely related strongly convex function
@@
This perturbed problem has minimum value at most $p_{\star}+\epsilon\frac{\|x_{\mathtt{OPT}}-x_{0}\|^{2}_{2}}{2D^{2}}$, and so any $\epsilon$-minimizer of $\tilde{f}_{0}$ is an $(1+\frac{\|x_{\mathtt{OPT}}-x_{0}\|^{2}_{2}}{2D^{2}})\epsilon$-minimizer for the original problem.
Note $\tilde{f}_{0}$ is $\epsilon/D^{2}$-strongly convex and since $m=0$, one can select $x_{\mathtt{SL}}=x_{\mathtt{OPT}}$. Moreover, if $f_{0}$ was $M$-Lipschitz continuous with a deterministic subgradient oracle, then as a sum of Lipschitz and smooth components, the perturbed objective $\tilde{f}_{0}$ satisfies (2.8) with $L_{0}^{2}=6M^{2}$ by Lemma 2.4. As a result, Theorem 3.2 ensures applying the subgradient method to $\tilde{f}_{0}$ with stepsize $\alpha_{k}=2D^{2}/\epsilon(k+2)$ has perturbed primal-dual gap converge at a rate $\frac{24M^{2}D^{2}}{\epsilon(T+1)}+O(1/T^{2})$. Similar perturbed primal-dual guarantees using a novel proof method were given by [27].
@@ 3.2 Proof of Primal-Dual Convergence Guarantees
This proof follows a standard analysis technique, directly expanding the definition of $R_{k+1}(y)$. First, suppose $x_{k}$ is feasible. Then
@@
where the first line uses that $\mathrm{prox}_{\alpha_{k},r}(y+\alpha_{k}n_{y})=y$, the second uses the nonexpansiveness of the proximal operator [34, Proposition 12.19], the third factors the norm squared and uses $\alpha_{k}=\lambda_{k}/w_{k}$, the fourth uses the strong convexity of $h_{y}$, and the fifth follows from definitions as $R_{k}(y)=\frac{w_{k}-\lambda_{k}\mu}{2}\mathbb{E}_{\xi}\|x_{k}-y\|_{2}^{2}$ and $\delta_{k}(y)=h_{y}(x_{k})-h_{y}(y)$. Then, applying the bound (2.8) gives the claim. Similarly, supposing $x_{k}$ is infeasible,
@@
$D_{k}:=\mathbb{E}_{\xi}\left[\left(\sum_{i<k:s(x_{i})=0}\lambda_{i}\right)p_{\star}-\inf M^{(k-1)}\right]\ .$ (3.8)
Note this relates to the dual gap as $D_{k}=\mathbb{E}_{\xi}\left[\left(\sum_{i<k:s(x_{i})=0}\lambda_{i}\right)\mathtt{dual\mbox{-}gap}_{k}\right]$.
@@
$D_{k+1}\leq D_{k}-\frac{\lambda_{k}}{2}\begin{cases}\left(2-L_{1}\alpha_{k}\right)\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}})-L_{0}^{2}\alpha_{k}&\text{if }x_{k}\text{ is feasible}\\ {\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}\left(2-L_{1}\alpha_{k}\right)\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}})+2\mathbb{E}_{\xi}f_{s(x_{k})}(x_{\mathtt{OPT}})-L_{0}^{2}\alpha_{k}}&\text{otherwise.}\end{cases}$
@@
Observe that one can rewrite $M^{(k)}(x)=Q_{1}(x)+Q_{2}(x)$ as the sum of two quadratics where $Q_{1}(x)=M^{(k-1)}(x)$ and $Q_{2}$ depends on whether $x_{k}$ is feasible. In the notation of Lemma 2.3, Lemmas 2.1 and 2.2 ensure $Q_{1}$ has $a_{1}=\inf M^{(k-1)}$, $b_{1}=\mu\sum_{i=0}^{k-1}\lambda_{i}$, and $z_{1}=x_{k}$, since $\bar{\beta}=0$ here. To determine $Q_{2}$, first suppose $x_{k}$ is feasible. Then we have
$Q_{2}(x)=\lambda_{k}\left(f_{0}(x_{k})+\langle g_{k},x-x_{k}\rangle+\frac{\mu}{2}\|x-x_{k}\|_{2}^{2}+r(x_{k+1})+\langle n_{k+1},x-x_{k+1}\rangle\right)\ .$
@@
$a_{2}$ $=\lambda_{k}\left(f_{0}(x_{k})+r(x_{k+1})+\alpha_{k}\langle n_{k+1},g_{k}+n_{k+1}\rangle-\frac{1}{2\mu}\|g_{k}+n_{k+1}\|_{2}^{2}\right)\ ,$
@@
From this, we conclude the lower bound
$\mathbb{E}_{\xi}\inf M^{(k)}$ $\geq\mathbb{E}_{\xi}\inf M^{(k-1)}+\lambda_{k}\mathbb{E}_{\xi}\Bigg{(}f_{0}(x_{k})+r(x_{\mathtt{OPT}})+\langle n_{x_{\mathtt{OPT}}},x_{k+1}-x_{\mathtt{OPT}}\rangle$
$\quad+\alpha_{k}\langle n_{k+1},g_{k}+n_{k+1}\rangle-\frac{1}{2\mu}\|g_{k}+n_{k+1}\|_{2}^{2}+\frac{\sum_{i=0}^{k-1}\lambda_{i}}{2\mu\sum_{i=0}^{k}\lambda_{i}}\left\|g_{k}+n_{k+1}\right\|_{2}^{2}\Bigg{)}$
$=\mathbb{E}_{\xi}\inf M^{(k-1)}+\lambda_{k}\mathbb{E}_{\xi}\Bigg{(}\delta_{k}(x_{\mathtt{OPT}})+p_{\star}+\langle n_{x_{\mathtt{OPT}}},-\alpha_{k}(g_{k}+n_{k+1})\rangle$
$\quad+\alpha_{k}\langle n_{k+1},g_{k}+n_{k+1}\rangle-\frac{\alpha_{k}}{2}\|g_{k}+n_{k+1}\|_{2}^{2}\Bigg{)}$
$=\mathbb{E}_{\xi}\inf M^{(k-1)}+\lambda_{k}\left(\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}})+p_{\star}-\frac{\alpha_{k}}{2}\mathbb{E}_{\xi}\|g_{k}+n_{x_{\mathtt{OPT}}}\|_{2}^{2}{\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}+\frac{\alpha_{k}}{2}\|n_{k+1}-n_{x_{\mathtt{OPT}}}\|_{2}^{2}}\right)$
${\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}\geq\mathbb{E}_{\xi}\inf M^{(k-1)}+\lambda_{k}\left(\mathbb{E}_{\xi}\delta_{k}(x_{\mathtt{OPT}})+p_{\star}-\frac{\alpha_{k}}{2}\mathbb{E}_{\xi}\|g_{k}+n_{x_{\mathtt{OPT}}}\|_{2}^{2}\right)}$
where the first inequality lower bounds $r(x_{k+1})$ by $r(x_{\mathtt{OPT}})+\langle n_{x_{\mathtt{OPT}}},x_{k+1}-x_{\mathtt{OPT}}\rangle$, the first equality applies the definitions of $\delta_{k}(x_{\mathtt{OPT}})$ in (2.7), $\alpha_{k}$ in (3.1) with $\bar{\beta}=0$, and that $x_{k+1}-x_{k}=-\alpha_{k}(g_{k}+n_{k+1})$, and the second equality combines and simplifies terms. In terms of $D_{k}=\mathbb{E}_{\xi}[\sum_{i<k:s(x_{i})=0}\lambda_{i}p_{\star}-\inf M^{(k-1)}]$, this gives the following recurrence
@@
Applying (2.8) gives the claim in this case. Now suppose $x_{k}$ is infeasible. Then, noting $Q_{2}$ minimizes at $x_{k}-\frac{g_{k}}{\mu}$, Lemma 2.2 ensures
@@
$\mathbb{E}_{\xi}\inf M^{(k-1)}+\lambda_{k}\mathbb{E}_{\xi}\left(f_{s(x_{k})}(x_{k})-\frac{1}{2\mu}\|g_{k}\|_{2}^{2}\right)+\frac{\lambda_{k}\sum_{i=0}^{k-1}\lambda_{i}}{{\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}2}\mu\sum_{i=0}^{k}\lambda_{i}}\mathbb{E}_{\xi}\left\|g_{k}\right\|_{2}^{2}$
@@
In terms of $D_{k}=\mathbb{E}_{\xi}[\sum_{i<k:s(x_{i})=0}\lambda_{i}p_{\star}-\mathbb{E}_{\xi}\inf M^{(k-1)}]$, this gives the recurrence
$D_{k+1}=D_{k}-\lambda_{k}\mathbb{E}_{\xi}\left[\delta_{k}(x_{\mathtt{OPT}})+f_{s(x_{k})}(x_{\mathtt{OPT}})\right]+\frac{\lambda_{k}\alpha_{k}}{2}\mathbb{E}_{\xi}\|g_{k}\|_{2}^{2}\ .$
Bounding $\mathbb{E}_{\xi}\|g_{k}\|_{2}^{2}$ by (2.8) gives the claim in this last case.
@@
$\mathbb{E}_{\xi}\|x_{T}-y\|^{2}_{2}$ $\leq\left(1+\frac{\max\{2,L_{1}/\mu-2\}L_{1}}{\mu}\right)^{T}\left(\|x_{0}-y\|^{2}_{2}+\frac{L_{0}^{2}}{{\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}L_{1}^{2}}}+\frac{L_{0}^{2}}{\mu\max\{2,L_{1}/\mu-2\}{\color[rgb]{0.0,0.0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.0,0.0}\pgfsys@color@gray@stroke{0.0}\pgfsys@color@gray@fill{0.0}L_{1}}}\right)\ .$
@@
The first inequality lower bounding $\delta_{k}(y)$ is immediate from (2.8). For the second inequality, note that if $x_{k}$ is feasible, convexity of $h_{y}$ ensures that $h_{y}(y)\geq h_{y}(x_{k})+\langle\mathbb{E}_{\xi_{k}}(g_{s(x_{k})}(x_{k};\xi_{k})+n_{y}),y-x_{k}\rangle$. If $x_{k}$ is infeasible, $f_{s(x_{k})}(y)\geq f_{s(x_{k})}(x_{k})+\langle\mathbb{E}_{\xi_{k}}g_{s(x_{k})}(x_{k};\xi_{k}),y-x_{k}\rangle$. In either case, Cauchy-Schwarz and Assumption C give the second inequality. Observe that if $\delta_{k}(y)<0$ , the first inequality above ensures $|\delta_{k}(y)|\leq\frac{L_{0}^{2}}{L_{1}}\leq L_{1}\|x_{k}-y\|_{2}^{2}+\frac{L_{0}^{2}}{L_{1}}$. Instead, if $\delta_{k}(y)\geq 0$, squaring the second inequality above ensures $\delta_{k}(y)^{2}\leq(L_{0}^{2}+L_{1}\delta_{k}(y))\|x_{k}-y\|^{2}_{2}$, which implies
$\delta_{k}(y)\leq\frac{L_{1}+\sqrt{L_{1}^{2}+4L_{0}^{2}/\|x_{k}-y\|_{2}^{2}}}{2}\|x_{k}-y\|_{2}^{2}\leq\frac{L_{0}^{2}}{L_{1}}+L_{1}\|x_{k}-y\|_{2}^{2},$
where the second inequality uses concavity to bound $\sqrt{a+b}\leq\sqrt{a}+\frac{b}{2\sqrt{a}}$, given $a>0$, completing the proposition’s first claim. To prove the proposition’s second claim, note
@@
$\leq\left(1+\frac{\max\{2,L_{1}/\mu-2\}L_{1}}{\mu}\right)\mathbb{E}_{\xi}\|x_{k}-y\|^{2}_{2}+\frac{\max\{2,L_{1}/\mu-2\}}{\mu}\frac{L_{0}^{2}}{L_{1}}+\frac{L_{0}^{2}}{\mu^{2}},$
where the first inequality uses Lemma 3.2 divided by $(\frac{\mu}{2}\sum_{i=0}^{k}\lambda_{i})$, the second inequality applies simple upper bounds, the third inequality uses our bound on $|\delta_{k}(y)|$, and the fourth uses that $0<\alpha_{k}\leq 1/\mu$ and its consequence $|2-L_{1}\alpha_{k}|\leq\max\{2,L_{1}/\mu-2\}$. From this, the proposition’s second claim follows as such recurrences of the form $a_{k+1}\leq b\cdot a_{k}+c$ satisfy $a_{T}\leq b^{T}a_{0}+c\frac{b^{T}-1}{b-1}\leq b^{T}(a_{0}+\frac{c}{b-1})$ , where $b>1$ and $c\geq 0$. ∎